# Beehive: A Flexible Network Stack for Direct-Attached Accelerators

### v3

@@ 3.4. Tile Chain Routing
In addition to NoC-level routing, Beehive has routing at the network packet level to determine the sequence of tiles that need to be chained together. We considered two routing methods: node-table routing, where each tile determines the correct next tile, and source routing, where the chain of tiles is completely determined when the first NoC message in the chain is created, such as when a packet is first received from the network. We chose to use node-table routing, because certain classes of traffic we want to support for our goal of interoperability require per-flow state or non-trivial protocol processing to fully determine the chain of tiles. Specifically, we consider routing for traffic that is either encrypted or is for layer 7.
@@
Encryption may obfuscate parts of packet payloads that are needed to fully route a packet, which would require the ingress tile to handle the decryption. An application request can span multiple packets. Which application tile should receive an RPC may depend on the RPC header or even the contents of the request. Further, the packets of one request can potentially be reordered or interleaved with other requests. To properly route these requests, an ingress tile would need to reassemble the stream, further complicating the implementation. In both cases, the ingress tile would need to implement significant, high-level protocol logic which is detrimental for modularity.
@@
(b) This tile assignment is able to avoid deadlock since the packet acquires NoC resources in order
Figure 5. A demonstration of how tile assignment affects deadlock. We can take advantage of protocol layer ordering, so a packet being processed always acquires NoC resources in the same order.
3.5. Deadlock
As with any NoC-based design, avoiding message-based deadlock must be a consideration. We note that NoC deadlock detection, avoidance, and recovery is a complex problem with a whole body of research behind it.
NoCs can deadlock in two ways: at the routing level and at the message passing level. To prevent routing-level deadlocks, we employ dimension-ordered routing (dally_deadlock_free_routing, ). Message passing deadlocks are a bigger concern in Beehive, because we enable each tile to route to any other tile at runtime. This means that our routing resources can get exhausted. The deadlock in Figure 5(a) is an example of this, in which the UDP RX tile must route east twice in one chain and it cannot route east a second time.
We apply resource acquisition ordering to solve this problem. Resource ordering can be imposed by taking advantage of the fact that protocol layers and services are composed in certain orders. For any individual packet, the path is determined at runtime, but we assume that all possible paths through the network stack for supported packet types are known when the network stack is compiled. As a simple example, consider the examples in Figure 5 of different topologies for the receive path of a UDP stack. Beehive’s NoC uses wormhole, dimension-ordered routing. The packet should be processed by Ethernet, IP, UDP, and then the application. With the tile layout in Figure 5(a), the route from the Ethernet to IP tile passes through the UDP tile’s router. As the UDP tile attempts to pass the packet along to the application, it must reacquire a NoC link it already used (5) and is thus deadlocked. If tiles are laid out as in Figure 5(b), no resources need to be reacquired, and the packet can be processed successfully.
@@ 3.5. Deadlock
Repeated protocol headers (e.g. two IP headers in the IP-in-IP protocol) break resource ordering. In Beehive, we choose to duplicate tiles (e.g. two IP RX tiles). If tiles are too expensive to duplicate, a potential solution is adding buffers to break dependencies (message_deadlock_recovery_avoidance_lankes_nocs, ; mdisha_song, ). These buffers give space for the NoC to drain into, freeing routing resources.
3.6. Control Plane Interfaces
For manageability, network operators need to be able to reconfigure protocol components from an external controller over a transport-layer connection. In Beehive, we choose to use a NoC as well for the control plane rather than a dedicated control bus. This is because control plane management can also benefit from a structured interconnect for scalability reasons.
First, for complex designs with a large number of components, it becomes costly to run dedicated, ad-hoc wires to every tile. Second, we want configuration to be over a reliable transport. This requires the control plane to use the transport layer, and a NoC enables this without physically coupling the component to the transport layer. This also enables us to add specific control plane management tiles to orchestrate state modifications. We describe a specific example in Section 4.5.
@@ 3.7. Application Interfaces
To support duplicated application tiles connected to the same transport layer, the NoC provides a convenient structure to multiplex between them in a scalable manner. The modularity provided by message passing on the NoC also allows an application to easily interface with any protocol in the network stack while reusing existing wires if, for example, an application wanted to switch from TCP to a custom reliable transport protocol. Finally, the standard interface of the NoC enables easy insertion of filters on the application’s NoC message, so network operators can enforce policies, such as dropping network traffic to or from non-whitelisted nodes. We describe the application NoC interface to our TCP layer in Section 4.4.
4. Implementation
To demonstrate the Beehive approach, we built a set of core protocol tiles, network functions, and applications. For protocols, we implement tiles for Ethernet, IPv4, UDP, and TCP. For network functions, we implement an IPinIP encapsulation layer and a NAT layer for network virtualization. For applications, we implement a Reed-Solomon encoder and an accelerator for a viewstamped replication node. These applications are described in more detail in Section 5.
We also describe our tooling that we developed to lower the effort required to maintain multiple designs and integrate new components. All of Beehive is implemented in standard SystemVerilog and was tested on an Alveo U200 communicating with standard CPU clients using a Linux or kernel-bypass network stack. We embed our Beehive prototype within Corundum (corundum_paper, ), an open-source 100 Gbps NIC, in the application slot to provide FPGA-specific infrastructure, such as the Ethernet MAC.
4.1. Network-on-chip (NoC)
We use the 2D mesh NoC from OpenPiton (openpiton_asplos, ) with some modifications. The NoC is wormhole-routed, uses dimension-ordered routing, and is full-duplex. We widen the NoC to 512 bits to match the width of the Xilinx MAC IP core and increase the flit width to 512 bits. The header flit format is inherited from OpenPiton. The maximum payload size for a NoC message is 256 MiB.
4.2. Protocol tiles
Protocols have one tile each for transmit and for receive processing. Each tile can be replicated if more throughput is needed for that element. Protocols are implemented as streaming components, so they begin to transmit the next NoC message as soon as possible rather than storing the entire NoC message before forwarding. Since each router has one input interface and one output interface for the tile, each side will utilize an entire router’s bandwidth if running at 100 Gbps. Since the packet-level protocol layers do not share state between their transmit and receive sides, this is a straightforward split.
@@ 4.2. Protocol tiles
Protocol tiles also have optional hash tables that use the 4 tuple as the key for load balancing to downstream replicated tiles. We set up initial packet-level routing within the tiles at compile time when we build the FPGA image. The hash table can be rewritten during runtime. Any packet that does not have an entry for a next hop (e.g.traffic with an unsupported protocol) is dropped to filter out unwanted traffic.
4.3. Buffer tiles
@@
In Beehive, we also have buffer tiles that hold large blocks of memory. In our current prototype, these buffers are large BRAMs, but the backing buffer could also be DRAM. These buffer tiles are accessible to any other tile in the system via NoC messages. This allows us to have shared buffers between tiles, so that multiple tiles can share state when needed.
4.4. TCP engine
@@
In order to evaluate how Beehive can support reliable transport, we prototype a TCP engine that implements server-side TCP. It can receive connection setup requests, generate sequence and ACK numbers, and support fast retransmit and window-based flow control (rfc_fast_retransmit, ). Currently, it does not support selective acknowledgments, initiating connections, or congestion control. Full TCP offload functionality has been demonstrated by previous work (limago, ) and could be integrated into Beehive.
@@
While the TCP engine has an RX router and a TX router like the other protocol tiles, the send and receive paths in TCP must share state. For example, the transmit path needs to know for which packets it has received acknowledgments. We choose to support sharing by running dedicated wires between the tiles. Every receive path only has one corresponding transmit path, so wires do not fan out. Although dedicated wires cause tiles to be closely coupled, the frequent NoC messages needed for state updates would lead to a practical coupling.
@@
On the completion of the 3-way handshake, the TCP engine sends a NoC message to notify an application tile based on the destination port for the connection. On the receive side, the TCP engine implements an interface that lets an application specify the size of the request it should be notified for with a NoC message. When enough data has arrived to satisfy that request, the TCP engine sends a notification message back to the application with the buffer address where the data requested has been stored. The application then retrieves the data from the buffer for processing before sending another message to the TCP stack when it has finished using the data.
@@
The transmit side is similar. The TCP engine implements an interface where the application can request space in its transmit buffer of a certain size. The TCP engine sends a notification when there is room in that buffer with the buffer address where the data should be stored. The application then copies the data into the buffer and notifies the TCP engine.
4.5. Network function tiles
We implement both IPinIP encapsulation and an IP NAT; the client-side TCP migration we use for our evaluation assumes a NAT instead of encapsulation. For both tiles, the control plane can dynamically update the table that maps virtual IPs to physical IPs. This mapping is changed when the client migrates.
@@ 4.5. Network function tiles
Taking advantage of Beehive’s control plane interconnect, we implement an internal controller as a separate tile that receives an RPC over TCP from an external controller. The internal controller then sends NoC messages to the IP encapsulation or NAT tiles with the information needed to update their tables. Finally, the internal controller sends a confirmation response to the external controller over the TCP connection.
4.6. Debugging and logging
@@
In Beehive, tiles may keep logs, and we provide a UDP-based protocol to readback logs. Each log is associated with a particular UDP port and exposes its own interface on the NoC for readback. The UDP receive layer is responsible for directing packets to the appropriate log interfaces. The log read interface keeps a small buffer for requests and drops requests when it is full. The client program reads out the log an entry at a time and resend requests for any entries for which it does not receive a response.
We also have tiles that log information about TCP packet headers to help provide more visibility into the FPGA’s execution. This log can later be replayed in a cycle-accurate simulation. We also found that having a cycle accurate trace is necessary for proper replay, because the TCP engine may behave differently depending on the timing of events (e.g. dropping different packets).
@@
These tiles have two interfaces available over the NoC. One is used to forward packets to and from the TCP engine. This path logs the header information with a cycle timestamp. The other interface allows the logs to be read out over the network in response to a request sent over UDP. Because the logging tiles are embedded within the fabric, they can record the exact timing that packets entered and exited the TCP engine. During simulation replay, Beehive’s modularity allows us to easily replace the logging tiles with an interface to our testing replay framework.
4.7. Tooling
@@
We developed a set of tools that lower the engineering effort to create new designs, such as generating portions of the Verilog (e.g. top-level wiring for NoCs) or performing compile-time deadlock analysis. The design configuration is passed to these tools via an XML file. The XML file contains the expected dimensions for the design as well as an element for each NoC tile endpoint. At minimum, this element contains tags specifying a name to use for the endpoint as well as its X and Y coordinates. It may also contain fields with information for generating the tables used for determining the correct next hops.
@@
Given the dimensions in the XML file, we generate declarations of all the top-level wires between tiles. We also generate the subset of the port connections for a tile that correspond to wires between NoC routers and connect the appropriate wires for the tile configuration. We choose not to generate the whole tile instantiation, because certain tiles need to maintain additional ports for I/O, such as the Ethernet MAC.
@@
The XML file also enables us to check whether the high-level topology of the NoC is sound. For example, we check two tiles have the same X and Y coordinates, and all NoC coordinates are within the expected dimensions of the design. Because a 2D mesh must be a rectangle, this also gives us the opportunity to automatically generate empty tiles that just contain a router. We also use the information about the NoC topology and next hops in the XML file to generate a resource dependency graph that we can analyze for cycles.
5. Integrating with Beehive
5.1. Erasure Coding
@@
Erasure codes such as Reed-Solomon (RS) are commonly used in distributed storage systems to achieve high resilience to disk failures with modest storage overhead (azureerasure, ; metaerasure, ; fasterasure, ). An RS encoder turns input data into data plus redundancy bits at a pre-set ratio, striped across storage servers. If some storage elements fail, the remaining blocks from the stripe can be combined with these extra blocks to regenerate the missing blocks.

### v5

@@ IV-D Tile chain routing
In addition to NoC-level routing, Beehive routes at the network packet level to determine the sequence of tiles that need to be chained together. We considered two routing methods: node-table routing, where each tile determines the correct next tile, and source routing, where the chain of tiles is completely determined when the first NoC message in the chain is created, such as when a packet is first received from the network. We use node-table routing, because certain classes of traffic we want to support for interoperability require per-flow state or non-trivial protocol processing to fully determine the chain of tiles.
@@
Specifically, we consider routing for traffic that is either encrypted or is for layer 7. Encryption may obfuscate parts of packet payloads that are needed to fully route a packet, which would require the ingress tile to handle the decryption. An application request can span multiple packets. Which application tile should receive an RPC may depend on the RPC header or even the contents of the request. Further, the packets of one request, which may not fit in the first packet, could be reordered or interleaved with other requests. To properly route such requests, an ingress tile would need to assemble or reorder the stream, further complicating the implementation. In both cases, the ingress tile would need to implement significant, high-level protocol logic which is detrimental for modularity.
@@
(b) This tile assignment is able to avoid deadlock since the packet acquires NoC resources in order.
Figure 5: An example of how tile assignment affects deadlock. Beehive takes advantage of protocol layer ordering, so a packet always acquires NoC resources in the same order.
IV-E Deadlock
As with any NoC-based design, avoiding message-based deadlock must be a consideration. We note that NoC deadlock detection, avoidance, and recovery is a complex problem with a whole body of research behind it [21, 46, 58, 3, 68].
NoCs can deadlock in two ways: at the routing level and at the message passing level. To prevent routing-level deadlocks, we employ dimension-ordered routing [21]. Message passing deadlocks are a bigger concern in Beehive, because any tile can route to any other tile at runtime. This means that our routing resources can get exhausted. The deadlock in Figure 5(a) is an example of this, in which the UDP RX tile must route east twice in one chain, and it cannot route east a second time.
We apply resource acquisition ordering to solve this problem. Resource ordering can be imposed by taking advantage of the fact that protocol layers and services are composed in certain orders. Although packet routing is dynamic, we assume that all possible paths through the network stack for supported packet types are known when the network stack is compiled. As a simple example, Figure 5 shows different topologies for the receive path of a UDP stack. Beehive’s NoC uses wormhole, dimension-ordered routing. The packet should be processed by Ethernet, IP, UDP, and then the application. With the tile layout in Figure 5(a), the route from the Ethernet to IP tile passes through the UDP tile’s router (2). As the UDP tile attempts to pass the packet along to the application (4), it must reacquire a NoC link that is still in use (5) and is thus deadlocked. If tiles are laid out as in Figure 5(b), no resources need to be reacquired, and the packet can be processed successfully.
@@ IV-E Deadlock
Repeated protocol headers (e.g. two IP headers in the IP-in-IP protocol) break resource ordering. In Beehive, we choose to duplicate tiles (e.g. two IP RX tiles). If tiles are too expensive to duplicate, a potential solution is adding buffers to break dependencies [46, 71]. These buffers give space for the NoC to drain into, freeing routing resources.
IV-F Control plane interfaces
For manageability, network operators need to be able to reconfigure protocol components from an external controller over a transport-layer connection. In Beehive, we choose to use an additional separate message-based, routed NoC for the control plane rather than a dedicated control bus. This is because control plane management also benefits from a structured interconnect for scalability reasons.
First, for complex designs with a large number of components, it becomes costly to run dedicated, ad-hoc wires to every tile. Second, we want configuration to be over a reliable transport. This requires the control plane to use the transport layer, and a NoC enables this without physically coupling the component to the transport layer. This also enables us to add specific control plane management tiles to orchestrate state modifications. We describe a specific example in Section V-E.
@@ IV-G Application interfaces
The NoC provides a convenient structure to multiplex between duplicated application tiles connected to the same transport layer in a scalable manner. The modularity provided by message passing on the NoC also allows an application to easily interface with any protocol in the network stack while reusing existing wires if, for example, we want to switch from TCP to a custom reliable transport protocol. Finally, the standardized NoC interface enables easy insertion of filters on the application’s NoC messages, so network operators can enforce policies, such as dropping network traffic to or from non-whitelisted nodes. We describe the application NoC interface to our TCP layer in Section V-D.
V Implementation
To demonstrate the Beehive approach, we built a set of core protocol tiles, network functions, and applications. For protocols, we implement tiles for Ethernet, IPv4, UDP, and TCP. For network functions, we implement an IPinIP encapsulation layer and a NAT layer for network virtualization. For applications, we implement a Reed-Solomon encoder and an accelerator for a viewstamped replication node. These applications are described in more detail in Section VI.
We also describe our tooling that we developed to lower the effort required to maintain multiple designs and integrate new components. All of Beehive is implemented in standard SystemVerilog and was tested on an Alveo U200 communicating with standard CPU clients using a Linux or kernel-bypass network stack. We embed our Beehive prototype within Corundum [31], an open-source 100 Gbps NIC, in the application slot to provide FPGA-specific infrastructure, such as the Ethernet MAC. Corundum does not provide any higher-level packet processing logic for Beehive.
V-A Network-on-chip (NoC)
We use the 2D mesh NoC from OpenPiton [7] with some modifications. The NoC is wormhole-routed, uses dimension-ordered routing, and is full-duplex. We widen the NoC from 64 bits to 512 bits to match the width of the Xilinx MAC IP core, so it has a maximum throughput of 128 Gbps when running at 250 MHz and increase the flit width to 512 bits. Because the NoC only relies on the top 64 bits of the first flit to do NoC routing, we are able to reuse the NoC without further modification by making the top 64 bits of our first header flit the same as the original NoC header. The maximum payload size for a NoC message is 256 MiB.
V-B Protocol tiles
Protocols are implemented as streaming components, so they begin to transmit the next NoC message as soon as possible rather than storing the entire NoC message before forwarding. This is done to reduce latency as header parsing can be overlapped with payload copying. This is especially important when chaining, because each layer of header adds an extra layer of parsing.
@@ V-B Protocol tiles
One of the more difficult aspects of removing the headers from network packets is that certain protocols (e.g. IP or TCP) allow headers to have options, so the headers are not a fixed width. This means removing a packet header often requires removing a variable number of bytes from the stream. We implement this by appending two lines of data and then using a shifter to remove the required amounts of bytes.
For a protocol, we place the receive and transmit engines in separate tiles. This is because they are streaming and each router has one input and one output interface, so one engine will utilize an entire router’s bandwidth if running at 100 Gbps. Since the packet-level protocol layers do not share state between their transmit and receive sides, this is a straightforward split. The exception to this is the TCP engine which we discuss further in Section V-D.
@@
Protocol tiles also have optional hash tables that use the 4 tuple as the key for load balancing to downstream replicated tiles. We set up initial packet-level routing within the tiles at compile time when we build the FPGA image. The hash table can be rewritten during runtime via the control plane described in Section IV-F. Any packet that does not have an entry for a next hop (e.g. traffic with an unsupported protocol) is dropped to filter out unwanted traffic.
V-C Buffer tiles
@@
In Beehive, we also have buffer tiles that hold large blocks of memory. In our current prototype, these buffers are large BRAMs, but the backing buffer can also be DRAM. These buffer tiles are accessible to any other tile in the system via NoC messages. This allows us to have shared buffers between tiles, so that multiple tiles can share state when needed.
V-D TCP engine
@@
To evaluate how Beehive can support reliable transport, we prototype a TCP engine that implements server-side TCP. It can receive connection setup requests, generate sequence and ACK numbers, and support fast retransmit and window-based flow control [10]. Currently, it does not support selective acknowledgments, initiating connections, or congestion control. Full TCP offload functionality has been demonstrated by previous work [66] and could be integrated into Beehive.
We split the TCP logic into receive and transmit engines. The receive engine is responsible for determining if received data is in order, calculating the next ACK, and processing ACKs for the transmitted data. The transmit engine is responsible for separating out buffers for sending and updating the sequence number for the transmitted stream.
We use two optimizations to handle state shared between the receive and transmit engines when they are both processing the same flow. We handle this in two ways. First, we divide flow state into two BRAMs by which engine writes the data to prevent write conflicts. Second, we take advantage of the asynchronous nature of the transmit and receive streams in TCP to tolerate slightly stale state and avoid bypassing state when the two engines are processing the same flow. For example, the transmit engine reads the current flow state with the ACK number for the received stream as $ACK\_RECV_{1}$ in cycle $n$. Meanwhile, the receive engine has processed a packet and updated the ACK number to $ACK\_RECV_{2}$ in cycle $n+1$. The transmit engine can still use $ACK\_RECV_{1}$ as long as it still uses all the other state it read in cycle $n$. Functionally, this is the same as if the received packet had been received slightly later and processed after the transmit engine had sent its packet, which is allowed due to the assumptions of TCP.
@@
While the TCP engine has an RX router and a TX router like the other protocol tiles, the send and receive paths in TCP must share state. For example, the transmit path needs to know for which packets it has received acknowledgments. We choose to support sharing by running dedicated wires between the tiles. Every receive path only has one corresponding transmit path, so wires do not fan out. We could implement state sharing over NoC messages, but the state is read and updated frequently, so the frequent NoC messages needed for state updates would encourage these tiles to be placed close to each other on the NoC anyway.
@@
On the completion of the 3-way handshake, the TCP engine sends a NoC message to notify an application tile based on the destination port for the connection. On the receive side, the TCP engine lets an application specify the size of the request it should be notified for with a NoC message. When enough data has arrived to satisfy that request, the TCP engine sends a notification message back to the application with the buffer address where the data requested has been stored. The application then retrieves the data from the buffer for processing before sending another message to the TCP stack when it has finished using the data.
@@
The TCP engine implements a similar interface for the transmit engine where the application can request space in its transmit buffer of a certain size. The TCP engine sends a notification when there is room in that buffer with the address where the data should be stored. The application then copies the data into the buffer and notifies the TCP engine.
V-E Network function tiles
@@
We implement both IPinIP encapsulation and an IP NAT. For both tiles, the control plane can dynamically update the table mapping virtual IPs to physical IPs, which occurs when the a client machine migrates. To change this mapping, we implement an internal controller as a separate tile that receives an RPC over TCP from an external controller. The internal controller utilizes the control NoC to send NoC messages to the IP encapsulation or NAT tiles with the information needed to update their tables. Finally, the internal controller sends a confirmation response to the external controller.
V-F Debugging and logging
@@
In Beehive, tiles may keep logs, and we provide UDP and TCP-based protocols to externally fetch logs. Each log is associated with a particular port and exposes an interface on the NoC to the network stack for readback. The layer 4 receive tiles are responsible for directing packets to the appropriate log interfaces. The log read interface keeps a small buffer for requests and drops requests when it is full. The client program reads out the log an entry at a time and resend requests for any entries for which it does not receive a response.
This logging ability was invaluable for debugging TCP when running on an FPGA. TCP is underspecified and the main verification is running against a common implementation, such as the Linux kernel [9], so we needed to run it on an FPGA to verify that it behaves as expected. The reduced visibility in this setting increases the difficulty of the already hard task of debugging a TCP implementation, due to the asynchronous and non-deterministic setting where certain bugs are dependent on the available bandwidth and loss events. As a result of the asynchrony, we need a cycle accurate trace for proper replay, because the TCP engine may behave differently depending on the timing of events (e.g. it may drop different packets). As a result of the bandwidth-dependence, we cannot rely on tcpdump to collect traces, because of the possibility different packets might be dropped by the engine versus tcpdump.
@@
We inserted tiles that log information about TCP packet headers into the processing between the TCP and IP layers. These tiles have two NoC interfaces: one is used to forward packets to and from the TCP engine and logs the header information with a cycle timestamp, the other interface allows the logs to be read out over the network in response to a request sent over UDP. Because the logging tiles are embedded within the fabric, they can record the exact timing and sequence of packets that entered and exited the TCP engine. Once this log is collected, we are able to replay the log in a cycle-accurate manner using the recorded timestamps by replacing the logging tiles with an interface to our trace replay framework.
V-G Tooling
@@
We developed a set of tools to lower the engineering effort to create new designs, such as generating portions of the Verilog (e.g. top-level wiring for NoCs) or performing compile-time deadlock analysis. The design configuration is passed to these tools via an XML file, which contains the design dimensions as well as an element for each NoC tile endpoint. At minimum, this element contains tags specifying a name to use for the endpoint as well as its X and Y coordinates. It may also contain fields with information for generating the tables used for determining the correct next hops.
@@
Given the dimensions in the XML file, we generate declarations of all the top-level wires between tiles. We also generate the subset of the port connections for each tile that correspond to wires between NoC routers and connect the appropriate wires for the tile configuration. We choose not to generate the whole tile instantiation, because certain tiles need to maintain additional ports for I/O, such as the Ethernet MAC.
@@
The XML file also enables us to check whether the high-level topology of the NoC is sound. For example, we check if two tiles have the same X and Y coordinates, and all NoC coordinates are within the expected dimensions of the design. Because a 2D mesh must be a rectangle, this also gives us the opportunity to automatically generate empty tiles that just contain a router, as in the bottom rightmost tile in the UDP stack shown in Figure 8(a). We also use information about the NoC topology and next hops in the XML file to generate a resource dependency graph that we analyze for cycles to ensure a deadlock-free design. Figure 6 is a visualization of the layout generated by the XML file for the consensus witness design in Section VI-B.
Figure 6: Beehive tile layout for Viewstamped Replication.
VI Integrating with Beehive
VI-A Erasure coding
@@
To demonstrate the benefits of Beehive for a throughput-oriented application, we integrate an accelerator for Reed-Solomon encoding. Erasure codes such as Reed-Solomon (RS) are commonly used in distributed storage systems to achieve high resilience to disk failures with modest storage overhead [34, 64, 41]. An RS encoder adds redundancy bits to input data at a pre-set ratio, striped across storage servers. If some storage elements fail, the remaining blocks from the stripe can be combined with these extra blocks to regenerate the missing blocks.